# Symplectic Lattices and GKP Codes - Simple Randomized Constructions from Cryptographic Lattices

### v1

@@ Abstract
We construct good GKP (Gottesman-Kitaev-Preskill) codes (in the sense of Conrad, Eisert and Seifert proposed) from standard short integer solution (SIS) lattices as well as from ring SIS and module SIS lattices, R-SIS and M-SIS lattices, respectively. These lattice are crucial for lattice-based cryptography. Our construction yields GKP codes with distance $\sqrt{n/\pi e}$. This compares favorably with the NTRU-based construction by Conrad et al. that achieves distance $\Omega(\sqrt{n/q}),$ with $n\leq q^{2}/0.28$. Unlike their codes, our codes do not have secret keys that can be used to speed-up the decoding. However, we present a simple decoding algorithm that, for many parameter choices, experimentally yields decoding results similar to the ones for NTRU-based codes. Using the R-SIS and M-SIS construction, our simple decoding algorithm runs in nearly linear time. Following Conrad, Eisert and Seifert’s work, our construction of GKP codes follows directly from an explicit, randomized construction of symplectic lattices with (up to constants $\approx 1$) minimal distance $(1/\sigma_{2n})^{1/2n}\approx\sqrt{\frac{n}{\pi e}}$, where $\sigma_{2n}$ is the volume of the $2n$-dimensional unit ball. Before this result, Buser and Sarnak gave a non-constructive proof for the existence of such symplectic lattices.
@@ 1 Introduction
Note that due to the constraint $n\leq q^{2}/0.28$, the distance bound $\Theta(\sqrt{n/q})$ scales like $\Theta(\sqrt[4]{n})$ rather than $\Theta(\sqrt{n})$ as required for a good family of GKP codes. Based on the so called Gaussian Heuristic (GH) (see [18, 23, 24]) and experiments in low dimensional experiments, Conrad et al. conjecture that variants of their NTRU construction lead to families of GKP codes that encode $n\log(\lambda),\lambda>1,$ logical qubits using $n$ bosonic modes and achieve code distance $\Delta\geq\sqrt{\frac{n}{\lambda\pi e}}$ (Conjecture 1 and 2 in [18]). They also provide arguments showing that replacing NTRU lattices by more general lattices based on $q$-ary symmetric matrices leads to GKP codes with distance $\Delta\geq\sqrt{\frac{n}{\lambda\pi e}}$ (see the introduction and Proposition 3 in [18]).
@@
    GKP codes based on M-SIS, and in particular R-SIS lattices, are closest to the NTRU construction in [18], hence they may shed more light onto those constructions,
@@
In addition to the blueprint from [18], we also use standard techniques to analyze the minimal distance of random SIS lattices. These analyses (for SIS and M-SIS lattices) are our main technical contribution. For SIS lattices, the analysis of the minimal distance closely follows [26]). However, as mentioned above, to obtain GKP codes we cannot choose arbitrary SIS lattices. Instead, their defining matrices must be symmetric. It turns out, that this poses no problem, as we prove that random symmetric matrices over finite fields share an important regularity property with general matrices: for any fixed vector v, its image under a uniformly random matrix (general or symmetric) is also uniformly distributed. For M-SIS lattices, the symmetry issue also exists and can be resolved as for SIS lattices. However, for M-SIS lattices a more difficult technical problem arises since we have to deal with equations over finite rings rather than finite fields. More precisely, for M-SIS lattices we deal with the ring $\mathbb{Z}_{q}[X](X^{n}+1)$, where $\mathbb{Z}_{q}=\mathbb{Z}/(q\mathbb{Z})$. Although we present a very general analysis, it leads to non-trivial results only for specific choices of $n$ and $q$. The easiest and probably most useful case occurs when $n$ is a power of $2$ and $q=3,5\mod 16$, in which case the polynomial $X^{n}+1$ factors over $\mathbb{Z}_{q}[X]$ into two irreducible polynomials with the same degree. Note that these choices for $n,q$ are also popular parameter choices in cryptography and they correspond to useful choices for the NTRU-based constructions in [18].
Using standard tools from lattice theory one can show that GKP codes from symplectic lattices can not achieve distance better than $\sqrt{2n/\lambda},$ where $n$ as before is the number of modes and $n\log(\lambda)$ is the number of encoded logical qubits. This bound follows easily from so called transference bounds [27] and similar bounds were also shown in [12]. This bound shows that our construction achieves, up to small constants, the best possible distance. The Gaussian Heuiristic (see [18]) indicates that the distance of our codes are even best possible, at least as long as one insists on achieving the distance with probability close to $1$.
@@
In Section 2 we provide all necessary concepts and results from lattice theory. The SIS-based construction of symplectic lattices is given in Section 3, the analysis is presented in Section 4, and the application to the construction of good GKP codes in Section 5. Here we also provide some well-known bounds on the invariants of symplectic lattices. In Section 6 we present our simple decoding algorithm. Section 7 summarizes our numerical findings for the SIS-based construction of GKP codes. In Sections 8 through 12 we show how to construct symplectic lattices from M-SIS lattices (Section 8), we analyze the construction (Section 9), apply it to the construction of GKP codes (Section 10), describe the adaptation of our decoding algorithm to the M-SIS-based construction (Section 11), and present and discuss the experimental results for the M-SIS-based construction (Section 12).
@@ 3 Constructing symplectic matrices from SIS lattices
So called modular lattices play an important role in modern lattice-based cryptography. In the most general case, such a lattice is defined via a matrix $\textbf{A}\in\mathbb{Z}^{n\times m},m>n,$ and an integer $q$. Then one sets
@@ 5 Main results
In this section we discuss several parameter choices in Theorem 12. Some of these choices lead to a constructive versions of a theorem by Buser and Sarnak [19] on the existence of symplectic lattices with bounded minmal distance, albeit with slightly worse constants than in [19]. Other choices lead to variants of a conjecture by Conrad, Eisert and Seifert [18] on the existence and construction of good GKP codes. In both cases, different parameter choices lead to different combinations of success probability and minimal lattice or code distance.
@@ 5.1 Symplectic lattices with large minimal distance
$(p/\sigma_{2n})^{1/2n}\geq\left(1-\frac{1}{2n^{3/4}}\right)\sqrt{\frac{n}{\pi e}}\quad\text{and $\sqrt{\frac{n}{2q}}\leq\frac{1}{2n^{3/4}}\sqrt{\frac{n}{\pi e}}$}.$
@@ 5.3 Upper bounds for minimal distance and covering radius and
In this section, we prove some upper bounds on the minimal distance of symlectic SIS lattices and GKP codes obtained from such lattices, as well as on the covering radius of symplectic SIS lattices. All bounds are either well-known or can be proven easily using techniques known from general SIS lattices. For the sake of completeness we include proofs.
@@ 6 Bounded distance decoding
Definition 19 (Problem $\textrm{BDD}_{R}$).
Given an basis B for lattice ${\cal L}={\cal L}(\textbf{B})$, a distance bound $R\in\mathbb{R}^{\geq 0}$ and a point $\textbf{t}\in\operatorname{span}({\cal L})$, such that $\operatorname{dist}(\textbf{t},{\cal L})\leq R$, find a lattice point $\textbf{v}\in{\cal L}$ with $\|\textbf{t}-\textbf{v}\|\leq R$. We denote this by $\textrm{BDD}({\cal L},\textbf{t},R)$.
Often, $R$ is expressed as a multiple of $\lambda_{1}({\cal L})$ or as a multiple of its covering radius $\rho({\cal L})$. Note, that if $R<\lambda_{1}({\cal L})/2$ then a solution to $\textrm{BDD}({\cal L},\textbf{t},R)$ is unique.
To solve the BDD problem for lattices $(\sqrt{\lambda}{\cal L})^{\times}=\frac{1}{\sqrt{\lambda}}{\cal L},$ where ${\cal L}=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ is symplectic, we propose the following simple algorithm BDD. The input to the algorithm is a symmetric matrix $\textbf{H}\in\mathbb{Z}^{n\times n}$, a prime $q$, an integer $\lambda$, together defining the lattice $\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$, and a target vector $v=(v_{1},\ldots,v_{n},v_{n+1},\ldots,v_{2n})$. It outputs an integer vector $c=(c_{1},\ldots,c_{n},c_{n+1},\ldots,c_{2n})$. Implicitly, the algorithm uses the basis $\frac{1}{\sqrt{\lambda q}}M(\textbf{H})$ for lattice $\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$ (see (4) and (5)).
$\textsc{Bdd}_{\text{triv}}$ on input $\textbf{H}\in\mathbb{Z}^{n\times n},q,\lambda\in\mathbb{N}$ and $\textbf{v}\in\mathbb{Z}^{2n}$: 1. For $i=1,\ldots,n$, set $c_{i}:=\lfloor\sqrt{\lambda q}v_{i}\rceil$. 2. Set $\textbf{c}^{\prime}:=(c_{1},\ldots,c_{n}),\textbf{w}=\textbf{v}^{\prime}-\textbf{c}^{\prime}\textbf{H}$, where $\textbf{v}^{\prime}:=(v_{n+1},\ldots,v_{2n})$, set $\textbf{w}=(w_{1},\ldots,w_{n})$. 3. For $i=n+1,\ldots,2n$, set $c_{i}:=\lfloor\sqrt{\lambda/q}w_{n-i}\rceil$. 4. Output $\textbf{c}=(c_{1},\ldots,c_{2n})$.
@@ 8.1 Basic definitions and results
$R:=\mathbb{Z}[X]/(\Phi(X))\quad\text{and}\quad R_{q}:=\mathbb{Z}_{q}[X]/(\Phi(X)).$
@@
The structure of rings $R=\mathbb{Z}[X]/(\Phi(X))$ and $R_{q}=\mathbb{Z}_{q}[X]/(\Phi(X))$ is determined by the factorization of the polynomial $\Phi(X)$ over $\mathbb{Q}$ modulo $q$. More precisely, if
$\Phi(X)=\prod_{i=1}^{m}p_{j}(X)\mod q$
for irreducible polynomials over $\mathbb{F}_{q}$ with $\deg(p_{j})=n_{j}$, then
$R_{q}\cong\bigotimes\mathbb{F}_{q}/(p_{j}(X))\cong\bigotimes_{i=1}^{m}\mathbb{F}_{q^{n_{j}}}.$ (11)
In the sequel, we only consider $\Phi(X)=X^{n}\pm 1$. Following [31], by $Q_{d}$ denote the $d$-th cyclotomic polynomial over some field $K$, that is,
$Q_{d}(X):=\prod_{\begin{subarray}{c}s=1\\ \gcd(s,d)=1\end{subarray}}^{d}(X-\zeta^{s}),$
where $\zeta$ denotes a primitive $d$-th root of unity over $K$. In our cases, $K=\mathbb{Q}$ or $K=\mathbb{F}_{q}$, where $q$ is prime and $\mathbb{F}_{q}$ denotes the field with $q$ elements. In any case, the degree $\deg(Q_{d}(X))$ is $\phi(d),$ where $\phi(\cdot)$ denotes Euler’s totient function. Then, over $\mathbb{Q}$ and over $\mathbb{F}_{q}$ and assuming $\gcd(n,q)=1$,
$X^{n}-1=\prod_{d\mid n}Q_{d}(X)\quad\text{and}\quad X^{n}+1=\prod_{\begin{subarray}{c}d\mid 2n\\ d\nmid n\end{subarray}}Q_{d}(X).$ (12)
In both cases, the coefficients of the polynomials $Q_{d}(X)$ are in $\mathbb{Q}$ and $\mathbb{F}_{q}$, respectively. Whereas in the case $K=\mathbb{Q}$ the polynomials $Q_{d}(X)$ are irreducible, this is, in general, not true for $\mathbb{F}_{q}$. Instead (see e.g. Theorem 2.47 in [31])
Theorem 21.
Let $K=\mathbb{F}_{q}$ and assume that $\gcd(d,q)=1$. Then $Q_{d}(X)$ factors into $\phi(d)/\ell$ distinct irreducible polynomial of degree $\ell$, where $\ell$ is the order of $q$ modulo $d$, i.e. $\ell$ is the smallest integer with $q^{\ell}=1\mod d$.
In applying this theorem below we use the Carmichael function $\lambda(\cdot)$, where for $n\in\mathbb{N}$, $\lambda(n)$ is the smallest integer $m$ such that
$\forall a\in\mathbb{Z}_{n}^{*}:a^{m}=1\mod n.$
[…]
Lemma 22.
Theorem 23.
If prime $q$ is not a primitive $\lambda$-root modulo $2n$, the theorem above can easily be adapted to reflect this, i. e. for each $d$ the number of terms and their degrees have to be adapted according to Theorem 21.

### v2

@@ Abstract
We construct good GKP (Gottesman-Kitaev-Preskill) codes (in the sense of Conrad, Eisert and Seifert proposed) from standard short integer solution (SIS) lattices as well as from ring SIS and module SIS lattices, R-SIS and M-SIS lattices, respectively. These lattices are crucial for lattice-based cryptography. Our construction yields GKP codes with distance $\sqrt{n/\lambda\pi e}$ for $\lambda^{n}$ logical dimensions. This compares favorably with the NTRU-based construction by Conrad et al. that achieves distance $\Omega(\sqrt{n/q}),$ with $n\leq q^{2}/0.28$. Unlike their codes, our codes do not have secret keys that can be used to speed-up the decoding. However, combining our results based on R-SIS with results by Stehlé and Steinfeld, we can prove Conjecture 2 from Conrad et al., namely with appropriate parameter choices, NTRU lattices yield GKP codes with distance $\sqrt{n/\pi e}$. Finally, we present a simple decoding algorithm that, for many parameter choices, experimentally yields decoding results similar to the ones for NTRU-based codes. Using the R-SIS and M-SIS construction, our simple decoding algorithm runs in nearly linear time. Following Conrad, Eisert and Seifert’s work, our construction of GKP codes follows directly from an explicit, randomized construction of symplectic lattices with (up to constants $\approx 1$) minimal distance $(1/\sigma_{2n})^{1/2n}\approx\sqrt{\frac{n}{\pi e}}$, where $\sigma_{2n}$ is the volume of the $2n$-dimensional unit ball. Before this result, Buser and Sarnak gave a non-constructive proof for the existence of such symplectic lattices.
@@ 1 Introduction
Note that due to the constraint $n\leq q^{2}/0.28$, the distance bound $\Theta(\sqrt{n/q})$ scales like $\Theta(\sqrt[4]{n})$ rather than $\Theta(\sqrt{n})$ as required for a good family of GKP codes. Based on the so called Gaussian Heuristic (GH) (see [11, 1, 31]) and experiments in low dimensional experiments, Conrad et al. conjecture that variants of their NTRU construction lead to families of GKP codes with $\lambda^{n}$ logical dimensions using $n$ bosonic modes and achieve code distance $\Delta\geq\sqrt{\frac{n}{\lambda\pi e}}$ (Conjecture 1 and 2 in [11]). They also provide arguments showing that replacing NTRU lattices by more general lattices based on $q$-ary symmetric matrices leads to GKP codes with distance $\Delta\geq\sqrt{\frac{n}{\lambda\pi e}}$ (see the introduction and Proposition 3 in [11]).
@@
    our main result on GKP codes based on R-SIS can be combined with results from Stehlé and Steinfeld [32] to prove Conjecture 2 by Conrad, Eisert and Seifert [11],
@@
In addition to the blueprint from [11], we also use standard techniques to analyze the minimal distance of random SIS lattices. These analyses (for SIS and M-SIS lattices) are our main technical contribution. For SIS lattices, the analysis of the minimal distance closely follows [26]. However, as mentioned above, to obtain GKP codes we cannot choose arbitrary SIS lattices. Instead, their defining matrices must be symmetric. It turns out that this poses no problem, as we prove that random symmetric matrices over finite fields share an important regularity property with general matrices: for any fixed vector v, its image under a uniformly random matrix (general or symmetric) is also uniformly distributed. For M-SIS lattices, the symmetry issue also exists and can be resolved as for SIS lattices. However, for M-SIS lattices a more difficult technical problem arises since we have to deal with equations over finite rings rather than finite fields. More precisely, for M-SIS lattices we deal with the ring $\mathbb{Z}_{q}[X]/(X^{n}+1)$, where $\mathbb{Z}_{q}=\mathbb{Z}/(q\mathbb{Z})$. Although we present a very general analysis, it leads to non-trivial results only for specific choices of $n$ and $q$. The easiest and probably most useful case occurs when $n$ is a power of $2$ and $q=3,5\mod 16$, in which case the polynomial $X^{n}+1$ factors over $\mathbb{Z}_{q}[X]$ into two irreducible polynomials with the same degree. Note that these choices for $n,q$ are also popular parameter choices in cryptography and they correspond to useful choices for the NTRU-based constructions in [11]. Our techniques are similar to the ones used by Stehlé and Steinfeld [32]. However, they focus on the $\ell_{\infty}$-norm, whereas we consider only the Euclidean norm. Moreover, since we are interested in exact constants we have to resolve additional problems stemming from the ring structure of $\mathbb{Z}_{q}[X]/(X^{n}+1)$. In particular, when proving the fundamental Lemma 27 we apply an averaging argument to deal with zero-divisors in $\mathbb{Z}_{q}[X]/(X^{n}+1)$. This argument is specific to the Euclidean norm.
Using standard tools from lattice theory one can show that GKP codes from symplectic lattices cannot achieve distance better than $\sqrt{2n/\lambda},$ where $n$ as before is the number of modes and $n\log(\lambda)$ is the number of encoded logical qubits. This bound follows easily from so-called transference bounds [4] and similar bounds were also shown in [10]. This bound shows that our construction achieves, up to small constants, the best possible distance. The Gaussian Heuristic (see [11]) indicates that the distance of our codes is even best possible, at least as long as one insists on achieving the distance with probability close to $1$.
@@
In Section 2 we provide all necessary concepts and results from lattice theory. The SIS-based construction of symplectic lattices is given in Section 3, the analysis is presented in Section 4, and the application to the construction of good GKP codes in Section 5. Here we also provide some well-known bounds on the invariants of symplectic lattices. In Section 6 we present our simple decoding algorithm. Section 7 summarizes our numerical findings for the SIS-based construction of GKP codes. In Sections 8 through 13 we show how to construct symplectic lattices from M-SIS lattices (Section 8), we analyze the construction (Section 9), apply it to the construction of GKP codes (Section 10), prove Conjecture 2 from [11] (Section 11), describe the adaptation of our decoding algorithm to the M-SIS-based construction (Section 12), and present and discuss the experimental results for the M-SIS-based construction (Section 13).
@@ 5 Main results
In this section we discuss several parameter choices in Theorem 12. Some of these choices lead to a constructive version of a theorem by Buser and Sarnak [7] on the existence of symplectic lattices with bounded minimal distance, albeit with slightly worse constants than in [7]. Other choices lead to variants of a conjecture by Conrad, Eisert and Seifert [11] on the existence and construction of good GKP codes. In both cases, different parameter choices lead to different combinations of success probability and minimal lattice or code distance.
@@ 5.1 Symplectic lattices with large minimal distance
$(p/\sigma_{2n})^{1/2n}\geq\left(1-\frac{1}{2n^{3/4}}\right)\sqrt{\frac{n}{\pi e}}\qquad\text{and $\sqrt{\frac{n}{2q}}\leq\frac{1}{2n^{3/4}}\sqrt{\frac{n}{\pi e}}$}.$
@@ 5.3 Upper bounds for minimal distance and covering radius and
In this section, we prove some upper bounds on the minimal distance of symplectic SIS lattices and GKP codes obtained from such lattices, as well as on the covering radius of symplectic SIS lattices. All bounds are either well-known or can be proven easily using techniques known from general SIS lattices. For the sake of completeness we include proofs.
@@ 6 Bounded distance decoding
Definition 19 (Problem $\textrm{BDDx}_{R}$).
Given a basis B for lattice ${\cal L}={\cal L}(\textbf{B})$, a distance bound $R\in\mathbb{R}^{\geq 0}$ and a point $\textbf{t}\in\operatorname{span}({\cal L})$, such that $\operatorname{dist}(\textbf{t},{\cal L})\leq R$, find a lattice point $\textbf{v}\in{\cal L}$ with $\|\textbf{t}-\textbf{v}\|\leq R$. We denote this by $\textrm{BDDx}({\cal L},\textbf{t},R)$.
Often, $R$ is expressed as a multiple of $\lambda_{1}({\cal L})$ or as a multiple of its covering radius $\rho({\cal L})$. Note, that if $R<\lambda_{1}({\cal L})/2$ then a solution to $\textrm{BDDx}({\cal L},\textbf{t},R)$ is unique.
To solve the BDD problem for lattices $(\sqrt{\lambda}{\cal L})^{\times}=\frac{1}{\sqrt{\lambda}}{\cal L},$ where ${\cal L}=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ is symplectic, we propose the following simple algorithm BDDx. The input to the algorithm is a symmetric matrix $\textbf{H}\in\mathbb{Z}^{n\times n}$, a prime $q$, an integer $\lambda$, together defining the lattice $\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$, and a target vector $v=(v_{1},\ldots,v_{n},v_{n+1},\ldots,v_{2n})$. It outputs an integer vector $c=(c_{1},\ldots,c_{n},c_{n+1},\ldots,c_{2n})$. Implicitly, the algorithm uses the basis $\frac{1}{\sqrt{\lambda q}}M(\textbf{H})$ for lattice $\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$ (see (4) and (5)).
$\textsc{Bdd}_{\text{triv}}$ on input $\textbf{H}\in\mathbb{Z}^{n\times n},q,\lambda\in\mathbb{N}$ and $\textbf{v}\in\mathbb{Z}^{2n}$: 1. For $i=1,\ldots,n$, set $c_{i}:=\lfloor\sqrt{\lambda q}v_{i}\rceil$. 2. Set $\textbf{c}^{\prime}:=(c_{1},\ldots,c_{n}),\textbf{w}=\textbf{v}^{\prime}-\textbf{c}^{\prime}\textbf{H}$, where $\textbf{v}^{\prime}:=(v_{n+1},\ldots,v_{2n})$, set $\textbf{w}=(w_{1},\ldots,w_{n})$. 3. For $i=n+1,\ldots,2n$, set $c_{i}:=\lfloor\sqrt{\lambda/q}w_{i-n}\rceil$. 4. Output $\textbf{c}=(c_{1},\ldots,c_{2n})$.
@@ 8.1 Basic definitions and results
$R:=\mathbb{Z}[X]/(\Phi(X))\qquad\text{and}\qquad R_{q}:=\mathbb{Z}_{q}[X]/(\Phi(X)).$
@@
In this paper we consider rings $R=\mathbb{Z}[X]/(X^{n}+1)$ and $\mathbb{Z}_{q}[X]/(X^{n}+1)$, where $n\in\mathbb{N}$ and $q$ a prime with $\gcd(n,q)=1$.The structure of rings $R=\mathbb{Z}[X]/(X^{n}+1)$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{n}+1)$ is determined by the factorization of the polynomial $X^{n}+1$ over $\mathbb{Q}$ and modulo $q$. More precisely, over $\mathbb{Q}$
$X^{n}+1=\prod_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}Q_{m}(X),$ (11)
where, following [22], $Q_{m}$ denotes the $m$-th cyclotomic polynomial over $\mathbb{Q}$. That is,
$Q_{m}(X):=\prod_{\begin{subarray}{c}s=1\\ \gcd(s,m)=1\end{subarray}}^{d}(X-\zeta^{s}),$
where $\zeta$ denotes a primitive $m$-th root of unity over $\mathbb{Q}$.
For $m\in\mathbb{N}$, set $m]:=\{0,\ldots,m\}$. Modulo $q$ the polynomials $Q_{m}$ are no longer irreducible, instead there exist integers $n_{m},d_{m}$ such that $Q_{m}$ factors into $n_{m}$ distinct irreducible factors $p_{m,j}$ each if degree $d_{m}$, where $d_{m}$ is the order of $q$ modulo $m$. (see e.g. Theorem 2.47 in [22]). Consequently, over the field $\mathbb{F}_{q}$, polynomial $X^{n}+1$ factors as
$X^{n}+1=\prod_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\prod_{j_{m}\in[n_{m}]}p_{m,j_{m}}(X),$ (12)
where the polynomials $p_{m,j_{m}}$ are pairwise distinct and irreducible. Hence
$R_{q}\cong\bigotimes_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\bigotimes_{j_{m}|in[n_{m}]}\mathbb{F}_{q}(X)/(p_{m,j_{m}})(X))\cong\bigotimes_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\left(\mathbb{F}_{q^{d_{m}}}\right)^{n_{m}}.$ (13)